OpenStudy (anonymous):

Find the quotient. Show all steps PLEASE. 4x^2 - y^2 / 2x + y

jimthompson5910 (jim_thompson5910):

\[\large \frac{4x^2 - y^2}{2x + y}\] \[\large \frac{(2x + y)(2x - y)}{2x + y}\] \[\large \frac{\cancel{(2x + y)}(2x - y)}{\cancel{2x + y}}\] \[\large 2x - y\] So \[\large \frac{4x^2 - y^2}{2x + y}=2x - y\]

OpenStudy (lalaly):

i love ur answers xD

jimthompson5910 (jim_thompson5910):

thx

OpenStudy (anonymous):

Jim is super fast!! Thanx Jim

jimthompson5910 (jim_thompson5910):

np glad to help